OpenStudy (anonymous):

Simplify 6xy(3xy)-2.

OpenStudy (anonymous):

See, 6xy(3xy) = ?? You have to just multiply 6xy with 3xy.. Can you do that??

OpenStudy (anonymous):

18x^2y^2-2=O

OpenStudy (anonymous):

multiply 6xy and 3xy, you can add the exponents of the variables to derive the right answer.

OpenStudy (anonymous):

What do you get? 6*3 = ??

OpenStudy (anonymous):

18

OpenStudy (anonymous):

I am not asking to you @HillDP

OpenStudy (anonymous):

aaay sorry i thought @iceywheniplease7 was asking. sorry @waterineyes .

OpenStudy (anonymous):

hi guys sorry i did know that it equaled 18 xD

OpenStudy (anonymous):

Okay. Now tell me : x*x = ??

OpenStudy (anonymous):

i got it, thnx